# Efficient Graph Optimization via Distance-Aware Graph Representation Learning

### v6

@@ A.1 Complete Proof of Theorem 1: Generalization Bound
Substituting the complexity bound completes the proof. ∎
A.2 Graph Perturbation and Error Decomposition
Let $\mathcal{E}$ denote the original edge set and $\mathcal{E}^{\prime}$ the modified edge set after applying DRTR. Assume the ground-truth underlying graph $\mathcal{G}^{*}$ satisfies a structural smoothness condition: for any edge $(u,v)\in\mathcal{E}^{*}$, we have $\|\mathbf{x}_{u}-\mathbf{x}_{v}\|_{2}^{2}\leq\delta$.
Now suppose the observed graph $\mathcal{G}$ contains a fraction $\epsilon$ of noisy edges. Let $\mathbf{z}_{v}$ and $\mathbf{z}_{v}^{*}$ denote the representations learned from $\mathcal{G}$ and $\mathcal{G}^{*}$, respectively.
We decompose the representation error as:
@@ A.2 Graph Perturbation and Error Decomposition
$\|\mathbf{z}_{v}-\mathbf{z}_{v}^{*}\|_{2}\leq\underbrace{\|\mathbf{z}_{v}-\mathbf{z}_{v}^{\text{DRTR}}\|_{2}}_{\text{Noise attenuation}}+\underbrace{\|\mathbf{z}_{v}^{\text{DRTR}}-\mathbf{z}_{v}^{*}\|_{2}}_{\text{Bias due to pruning}}.$
A.3 Bounded Noise Reduction via Distance Thresholding
We now formalize how DRTR filters out noisy edges using a distance threshold $\alpha$.
Proposition 1 (Noise Attenuation).
Suppose node features are $\ell$-Lipschitz over the clean graph $\mathcal{G}^{*}$. Let DRTR prune any edge $(u,v)$ where $\|\mathbf{x}_{u}-\mathbf{x}_{v}\|_{2}^{2}>\alpha$. Then the variance of $\mathbf{z}_{v}$ satisfies:
$\mathrm{Var}[\mathbf{z}_{v}^{\text{DRTR}}]\leq\mathrm{Var}[\mathbf{z}_{v}]-\epsilon\cdot C(\alpha,\ell),$
for some function $C(\alpha,\ell)>0$ increasing in $\alpha$.
Proof of Noise Attenuation.
The proof follows from the fact that noisy edges contribute additional variance to the representation. By removing edges with distance $\alpha$, we eliminate the variance contribution from these noisy connections.
Let $\mathcal{E}_{\text{noisy}}$ denote the set of noisy edges. The variance reduction is:
$\mathrm{Var}[\mathbf{z}_{v}]-\mathrm{Var}[\mathbf{z}_{v}^{\text{DRTR}}]=\sum_{(u,v)\in\mathcal{E}_{\text{noisy}}}\alpha_{vu}^{2}\cdot\mathrm{Var}[\mathbf{W}\mathbf{x}_{u}]\geq\epsilon\cdot C(\alpha,\ell),$
@@ A.3 Bounded Noise Reduction via Distance Thresholding
where $C(\alpha,\ell)=\min_{(u,v)\in\mathcal{E}_{\text{noisy}}}\alpha_{vu}^{2}\cdot\mathrm{Var}[\mathbf{W}\mathbf{x}_{u}]$ is positive and increasing in $\alpha$. ∎
A.4 Optimization Analysis and Convergence Guarantees
A.4.1 Objective Function Decomposition
The DRTR optimization problem can be formulated as:
$\min_{\theta}\mathcal{L}(\theta)=\mathcal{L}_{\text{classification}}(\theta)+\lambda_{1}\mathcal{L}_{\text{DR}}(\theta)+\lambda_{2}\mathcal{L}_{\text{TR}}(\theta)+\lambda_{3}\mathcal{L}_{\text{regularization}}(\theta),$
@@ A.4.1 Objective Function Decomposition
$\mathcal{L}_{\text{TR}}(\theta)$ $=\sum_{(v,u)\in\mathcal{E}_{\text{candidate}}}\left[s_{vu}\log P(\text{add edge }(v,u))+(1-s_{vu})\log(1-P(\text{add edge }(v,u)))\right]$ (21)
@@
A.4.2 Gradient Analysis
The gradient of the objective function with respect to parameters $\theta$ is:
$\nabla\mathcal{L}(\theta)=\nabla\mathcal{L}_{\text{classification}}(\theta)+\lambda_{1}\nabla\mathcal{L}_{\text{DR}}(\theta)+\lambda_{2}\nabla\mathcal{L}_{\text{TR}}(\theta)+\lambda_{3}\nabla\mathcal{L}_{\text{regularization}}(\theta).$
For the distance recomputation component:
$\frac{\partial\mathcal{L}_{\text{DR}}}{\partial\mathbf{W}^{(k)}}=\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{N}^{(k)}(v)}\mathbb{I}[d_{vu}^{(k)}>\alpha_{k}]\cdot\frac{\partial d_{vu}^{(k)}}{\partial\mathbf{W}^{(k)}},$
where $\mathbb{I}[\cdot]$ is the indicator function.
For the topology reconstruction component:
[…]
A.4.3 Convergence Rate Analysis
Theorem 5 (Convergence Rate of DRTR).
Under the following assumptions:
  • The objective function $\mathcal{L}(\theta)$ is $L$-smooth (i.e., $\|\nabla\mathcal{L}(\theta_{1})-\nabla\mathcal{L}(\theta_{2})\|_{2}\leq L\|\theta_{1}-\theta_{2}\|_{2}$)
  • The gradients are bounded: $\mathbb{E}[\|\nabla\mathcal{L}(\theta)\|_{2}^{2}]\leq G^{2}$
  • The learning rate satisfies $\eta_{t}=\frac{\eta_{0}}{\sqrt{t}}$ for some $\eta_{0}>0$
Then, the DRTR algorithm achieves:
@@ A.4.3 Convergence Rate Analysis
Convergence Proof.
The proof follows standard stochastic gradient descent analysis with adaptive learning rates.
@@
$\mathcal{L}(\theta_{t+1})\leq\mathcal{L}(\theta_{t})+\langle\nabla\mathcal{L}(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\frac{L}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}.$
@@
$\theta_{t+1}=\theta_{t}-\eta_{t}\nabla\mathcal{L}(\theta_{t}).$
@@
$\mathbb{E}[\mathcal{L}(\theta_{t+1})]\leq\mathcal{L}(\theta_{t})-\eta_{t}\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}+\frac{L\eta_{t}^{2}G^{2}}{2}.$
@@
$\sum_{t=1}^{T}\eta_{t}\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}\leq\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+\frac{LG^{2}}{2}\sum_{t=1}^{T}\eta_{t}^{2}.$
@@
$\sum_{t=1}^{T}\eta_{t}=\eta_{0}\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\geq\eta_{0}\sqrt{T},$
@@
$\min_{t=1,\ldots,T}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}]\leq\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+\frac{LG^{2}\eta_{0}^{2}\log T}{2}}{\eta_{0}\sqrt{T}}.$
This completes the convergence rate analysis. ∎
A.5 Computational Complexity Analysis
A.5.1 Time Complexity
The computational complexity of DRTR can be analyzed component-wise:
Distance Recomputator (DR):
  • Distance computation: $\mathcal{O}(K\cdot|\mathcal{E}|\cdot d)$
  • Thresholding: $\mathcal{O}(K\cdot|\mathcal{V}|\cdot\log|\mathcal{V}|)$
  • Total DR complexity: $\mathcal{O}(K\cdot|\mathcal{E}|\cdot d)$
Topology Reconstructor (TR):
  • Similarity computation: $\mathcal{O}(|\mathcal{V}|^{2}\cdot d)$
  • Edge addition: $\mathcal{O}(|\mathcal{E}_{\text{candidate}}|)$
  • Total TR complexity: $\mathcal{O}(|\mathcal{V}|^{2}\cdot d)$
Multi-hop Aggregation:
  • Attention computation: $\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2})$
  • Message passing: $\mathcal{O}(K\cdot|\mathcal{V}|\cdot d^{2})$
  • Total aggregation complexity: $\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2})$
Overall Complexity:
$\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2}+|\mathcal{V}|^{2}\cdot d)$
A.5.2 Space Complexity
The space complexity is dominated by:
  • Feature matrices: $\mathcal{O}(|\mathcal{V}|\cdot d)$
  • Attention matrices: $\mathcal{O}(K\cdot|\mathcal{E}|)$
  • Distance matrices: $\mathcal{O}(K\cdot|\mathcal{E}|)$
  • Total space complexity: $\mathcal{O}(|\mathcal{V}|\cdot d+K\cdot|\mathcal{E}|)$
A.6 Stability Analysis
Theorem 6 (Stability of DRTR).

### v7

@@ A.1 Complete Proof of Theorem 1 (Main Text): Generalization Bound
Substituting the complexity bound from Step 4 and the effective degree reduction from Step 2 completes the proof.
A.2 Auxiliary Lemmas for Generalization Bound
We first establish several key lemmas that will be used in the main proof. These lemmas control the error propagation through the multi-hop aggregation structure and establish the properties of our learning rate schedule.
Lemma A.3 (Properties of Learning Rate Weights).
For the learning rate $\alpha_{t}=\frac{H+1}{H+t}$ and the weights $\alpha_{i}^{t}$ defined as:
$\alpha_{0}^{t}=\prod_{j=1}^{t}(1-\alpha_{j}),\quad\alpha_{i}^{t}=\alpha_{i}\prod_{j=i+1}^{t}(1-\alpha_{j}),$ (25)
the following properties hold:
  1. (a)
    $\frac{1}{\sqrt{t}}\leq\sum_{i=1}^{t}\frac{\alpha_{i}^{t}}{\sqrt{i}}\leq\frac{2}{\sqrt{t}}$ for every $t\geq 1$.
  2. (b)
    $\max_{i\in[t]}\alpha_{i}^{t}\leq\frac{2H}{t}$ and $\sum_{i=1}^{t}(\alpha_{i}^{t})^{2}\leq\frac{2H}{t}$ for every $t\geq 1$.
  3. (c)
    $\sum_{t=i}^{\infty}\alpha_{i}^{t}=1+\frac{1}{H}$ for every $i\geq 1$.
Proof A.4 (Proof of Lemma A.0).
The proof follows from direct manipulation of the definition of $\alpha_{t}$ and $\alpha_{i}^{t}$.
Part (a): The proof is by induction on $t$. For $t=1$, we have $\sum_{i=1}^{1}\frac{\alpha_{i}^{1}}{\sqrt{i}}=\alpha_{1}=\frac{H+1}{H+1}=1$, so the statement holds. For $t\geq 2$, by the relationship $\alpha_{i}^{t}=(1-\alpha_{t})\alpha_{i}^{t-1}$ for $i=1,\ldots,t-1$, we have:
$\sum_{i=1}^{t}\frac{\alpha_{i}^{t}}{\sqrt{i}}=\frac{\alpha_{t}}{\sqrt{t}}+(1-\alpha_{t})\sum_{i=1}^{t-1}\frac{\alpha_{i}^{t-1}}{\sqrt{i}}.$ (26)
By induction and the fact that $\alpha_{t}=\frac{H+1}{H+t}$, the bounds follow.
Part (b): We have:
$\alpha_{i}^{t}=\frac{H+1}{i+H}\cdot\frac{i+1}{i+1+H}\cdots\frac{t-1}{t-1+H}\leq\frac{H+1}{i+H}\cdot\frac{t}{t+H}\leq\frac{2H}{t},$ (27)
where the last inequality uses $H\geq 1$. The second inequality follows since $\sum_{i=1}^{t}(\alpha_{i}^{t})^{2}\leq[\max_{i\in[t]}\alpha_{i}^{t}]\cdot\sum_{i=1}^{t}\alpha_{i}^{t}=[\max_{i\in[t]}\alpha_{i}^{t}]\cdot 1$.
Part (c): Using the identity for telescoping products:
$\sum_{t=i}^{\infty}\alpha_{i}^{t}=\frac{H+1}{i+H}\cdot\left(1+\frac{i}{i+1+H}+\frac{i}{i+1+H}\cdot\frac{i+1}{i+2+H}+\cdots\right)=\frac{H+1}{H}=1+\frac{1}{H}.$ (28)
This completes the proof.
Lemma A.5 (Recursion on Representation Error).
For any node $v\in\mathcal{V}$ and hop $k\in[K]$, let $t=N_{k}(v)$ and suppose node $v$ was processed at hop $k$ in iterations $t_{1},\ldots,t_{t}<t$. Then the representation error satisfies:
$(\mathbf{z}_{v}^{t}-\mathbf{z}_{v}^{*})=\sum_{i=1}^{t}\alpha_{i}^{t}\cdot(\mathbf{z}_{u_{i}}^{t_{i}}-\mathbf{z}_{u_{i}}^{*})+\beta_{t},$ (29)
where $\alpha_{i}^{t}$ are the learning rate weights, $u_{i}$ are the neighbors used in aggregation, and $\beta_{t}$ is a confidence bonus term.
Proof A.6 (Proof of Lemma A.1).
The proof follows from the update equation (4.1) and the definition of $\alpha_{i}^{t}$ in (4.2). By expanding the recursive update:
$\mathbf{z}_{v}^{t}=\alpha_{0}^{t}\mathbf{z}_{v}^{0}+\sum_{i=1}^{t}\alpha_{i}^{t}\left[\mathbf{W}^{(k)}\mathbf{x}_{u_{i}}+\mathbf{b}_{i}\right],$ (30)
where $\mathbf{b}_{i}$ is the confidence bonus. Subtracting the optimal representation $\mathbf{z}_{v}^{*}$ and using the Bellman-like optimality condition completes the proof.
Lemma A.7 (Bound on Representation Error).
There exists an absolute constant $c>0$ such that, for any $p\in(0,1)$, with probability at least 1p1-p, the following holds simultaneously for all $(v,k,t)\in\mathcal{V}\times[K]\times[T]$:
$0\leq(\mathbf{z}_{v}^{t}-\mathbf{z}_{v}^{*})^{\top}\mathbf{z}_{v}^{*}\leq\alpha_{0}^{t}H+\sum_{i=1}^{t}\alpha_{i}^{t}(\mathbf{z}_{u_{i}}^{t_{i}}-\mathbf{z}_{u_{i}}^{*})^{\top}\mathbf{z}_{u_{i}}^{*}+\beta_{t},$ (31)
where $\beta_{t}=2\sum_{i=1}^{t}\alpha_{i}^{t}b_{i}\leq 4c\sqrt{H^{3}\iota/t}$ and $\iota=\log(SAT/p)$.
Proof A.8 (Proof of Lemma A.2).
The proof uses Azuma-Hoeffding concentration inequality and the fact that the martingale difference sequence has bounded variance. For each fixed $(v,k)$, the sequence $\{(\mathbf{z}_{v}^{t}-\mathbf{z}_{v}^{*})^{\top}\mathbf{z}_{v}^{*}\}_{t=1}^{T}$ forms a martingale with respect to the filtration generated by previous updates. By Azuma-Hoeffding:
$\mathbb{P}\left[\sum_{i=1}^{t}\alpha_{i}^{t}\xi_{i}>c\sqrt{H^{3}\iota/t}\right]\leq\frac{p}{SAT},$ (32)
where $\xi_{i}$ are the martingale differences. A union bound over all $(v,k,t)$ completes the proof.
A.3 Graph Perturbation and Error Decomposition
For ground-truth graph $\mathcal{G}^{*}$ with structural smoothness ($\|\mathbf{x}_{u}-\mathbf{x}_{v}\|_{2}^{2}\leq\delta$ for $(u,v)\in\mathcal{E}^{*}$) and observed graph $\mathcal{G}$ with $\epsilon$-fraction noisy edges, the representation error decomposes as:
@@ A.3 Graph Perturbation and Error Decomposition
$\|\mathbf{z}_{v}-\mathbf{z}_{v}^{*}\|_{2}\leq\underbrace{\|\mathbf{z}_{v}-\mathbf{z}_{v}^{\text{DRTR}}\|_{2}}_{\text{Noise attenuation}}+\underbrace{\|\mathbf{z}_{v}^{\text{DRTR}}-\mathbf{z}_{v}^{*}\|_{2}}_{\text{Bias due to pruning}}.$ (33)
A.4 Bounded Noise Reduction via Distance Thresholding
Proposition A.9 (Noise Attenuation).
For $\ell$-Lipschitz features over clean graph $\mathcal{G}^{*}$, pruning edges with $\|\mathbf{x}_{u}-\mathbf{x}_{v}\|_{2}^{2}>\alpha$ yields:
$\mathrm{Var}[\mathbf{z}_{v}^{\text{DRTR}}]\leq\mathrm{Var}[\mathbf{z}_{v}]-\epsilon\cdot C(\alpha,\ell),$ (34)
@@ A.4 Bounded Noise Reduction via Distance Thresholding
where $C(\alpha,\ell)=\min_{(u,v)\in\mathcal{E}_{\text{noisy}}}\alpha_{vu}^{2}\cdot\mathrm{Var}[\mathbf{W}\mathbf{x}_{u}]>0$ increases with $\alpha$.
A.5 Optimization Analysis and Convergence Guarantees
A.5.1 Objective Function Decomposition
The DRTR optimization problem:
$\min_{\theta}\mathcal{L}(\theta)=\mathcal{L}_{\text{classification}}(\theta)+\lambda_{1}\mathcal{L}_{\text{DR}}(\theta)+\lambda_{2}\mathcal{L}_{\text{TR}}(\theta)+\lambda_{3}\mathcal{L}_{\text{regularization}}(\theta),$ (35)
@@ A.5.1 Objective Function Decomposition
$\mathcal{L}_{\text{TR}}(\theta)$ $=\sum_{(v,u)\in\mathcal{E}_{\text{candidate}}}\max(0,\beta-\text{sim}_{vu})\cdot P(\text{add edge }(v,u))$ (38)
@@
A.5.2 Gradient Analysis
The gradient:
$\nabla\mathcal{L}(\theta)=\nabla\mathcal{L}_{\text{classification}}(\theta)+\lambda_{1}\nabla\mathcal{L}_{\text{DR}}(\theta)+\lambda_{2}\nabla\mathcal{L}_{\text{TR}}(\theta)+\lambda_{3}\nabla\mathcal{L}_{\text{regularization}}(\theta),$ (40)
with component gradients:
$\frac{\partial\mathcal{L}_{\text{DR}}}{\partial\mathbf{W}^{(k)}}$ $=\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{N}^{(k)}(v)}\mathbb{I}[d_{vu}^{(k)}>\alpha_{k}]\cdot\frac{\partial d_{vu}^{(k)}}{\partial\mathbf{W}^{(k)}},$ (41)
$\frac{\partial\mathcal{L}_{\text{TR}}}{\partial\omega_{i}}$ $=\sum_{(v,u)\in\mathcal{E}_{\text{candidate}}}\mathbb{I}[\beta-\text{sim}_{vu}>0]\cdot P(\text{add edge }(v,u))\cdot\frac{\partial\text{sim}_{vu}}{\partial\omega_{i}}.$ (42)
A.5.3 Convergence Rate Analysis
Theorem A.10 (Convergence Rate of DRTR).
Under the assumptions that the objective function $\mathcal{L}(\theta)$ is $L$-smooth (i.e., $\|\nabla\mathcal{L}(\theta_{1})-\nabla\mathcal{L}(\theta_{2})\|_{2}\leq L\|\theta_{1}-\theta_{2}\|_{2}$), the gradients are bounded such that $\mathbb{E}[\|\nabla\mathcal{L}(\theta)\|_{2}^{2}]\leq G^{2}$, and the learning rate satisfies $\eta_{t}=\frac{\eta_{0}}{\sqrt{t}}$ for some $\eta_{0}>0$, the DRTR algorithm achieves:
@@ A.5.3 Convergence Rate Analysis
Proof A.11 (Convergence Proof).
Since DRTR preserves $L$-smoothness and bounded gradients, standard SGD analysis applies. The proof proceeds in several steps.
@@
$\mathcal{L}(\theta_{t+1})\leq\mathcal{L}(\theta_{t})+\langle\nabla\mathcal{L}(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\frac{L}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}.$ (43)
@@
$\theta_{t+1}=\theta_{t}-\eta_{t}\nabla\mathcal{L}(\theta_{t}).$ (44)
@@
$\mathbb{E}[\mathcal{L}(\theta_{t+1})]\leq\mathcal{L}(\theta_{t})-\eta_{t}\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}+\frac{L\eta_{t}^{2}G^{2}}{2},$ (45)
where $G^{2}=\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}]$ is the bound on gradient variance.
@@
$\sum_{t=1}^{T}\eta_{t}\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}\leq\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+\frac{LG^{2}}{2}\sum_{t=1}^{T}\eta_{t}^{2},$ (46)
where $\mathcal{L}^{*}$ is the optimal objective value.
@@
$\sum_{t=1}^{T}\eta_{t}=\eta_{0}\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\geq\eta_{0}\sqrt{T},$ (47)
@@
$\min_{t=1,\ldots,T}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}]\leq\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+\frac{LG^{2}\eta_{0}^{2}\log T}{2}}{\eta_{0}\sqrt{T}}=\mathcal{O}\left(\frac{1}{\sqrt{T}}\right).$ (49)
This completes the convergence rate analysis.
A.6 Computational Optimization Analysis
A.6.1 Time Complexity Optimization
The computational complexity of DRTR is optimized through component-wise analysis. The Distance Recomputator (DR) module performs distance computation:
$T_{\text{DR}}=K\cdot\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{N}^{(k)}(v)}\mathcal{O}(d)=\mathcal{O}(K\cdot|\mathcal{E}|\cdot d),$ (50)
with thresholding operations:
$T_{\text{threshold}}=K\cdot|\mathcal{V}|\cdot\mathcal{O}(\log|\mathcal{V}|)=\mathcal{O}(K\cdot|\mathcal{V}|\cdot\log|\mathcal{V}|),$ (51)
yielding total DR complexity $T_{\text{DR}}^{\text{total}}=\mathcal{O}(K\cdot|\mathcal{E}|\cdot d)$ (thresholding is dominated by distance computation).
The Topology Reconstructor (TR) module uses approximate $k$-nearest neighbor search, reducing candidate edges from $\mathcal{O}(|\mathcal{V}|^{2})$ to $\mathcal{O}(|\mathcal{V}|\cdot k)$:
$T_{\text{TR}}=\mathcal{O}(|\mathcal{V}|\cdot k\cdot\log|\mathcal{V}|)+\mathcal{O}(|\mathcal{V}|\cdot k\cdot d)=\mathcal{O}(|\mathcal{V}|\cdot k\cdot d),$ (52)
where $k\ll|\mathcal{V}|$ (typically $k=50$).
The multi-hop aggregation component involves attention computation:
$T_{\text{attention}}=K\cdot\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{N}^{(k)}(v)}\mathcal{O}(d^{2})=\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2}),$ (53)
and message passing:
$T_{\text{message}}=K\cdot|\mathcal{V}|\cdot\mathcal{O}(d^{2})=\mathcal{O}(K\cdot|\mathcal{V}|\cdot d^{2}).$ (54)
Combining these components, the overall time complexity is:
$T_{\text{DRTR}}=T_{\text{attention}}+T_{\text{message}}+T_{\text{DR}}^{\text{total}}+T_{\text{TR}}=\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2}+|\mathcal{V}|\cdot k\cdot d),$ (55)
which scales linearly with graph size $|\mathcal{V}|$ and edge count $|\mathcal{E}|$, making DRTR suitable for large-scale graphs.
A.6.2 Computational Speedup Analysis
Compared to naive multi-hop aggregation without optimization, DRTR achieves computational speedup through:
$\text{Speedup}=\frac{T_{\text{naive}}}{T_{\text{DRTR}}}=\frac{\mathcal{O}(K\cdot|\mathcal{V}|^{2}\cdot d^{2})}{\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2}+|\mathcal{V}|\cdot k\cdot d)}\approx\frac{|\mathcal{V}|^{2}}{|\mathcal{E}|+|\mathcal{V}|\cdot k/d},$ (56)
where $|\mathcal{E}|\ll|\mathcal{V}|^{2}$ for sparse graphs, yielding significant speedup. The effective neighborhood reduction factor is:
$\rho=\frac{d_{\text{eff}}}{d_{\text{original}}}=\frac{\sum_{v}|\mathcal{N}_{\text{eff}}(v)|}{\sum_{v}|\mathcal{N}_{\text{original}}(v)|}\ll 1,$ (57)
leading to computational reduction proportional to $\rho$.
A.6.3 Space Complexity
The space complexity is:
$S_{\text{DRTR}}=\mathcal{O}(|\mathcal{V}|\cdot d)+\mathcal{O}(K\cdot|\mathcal{E}|)+\mathcal{O}(K\cdot|\mathcal{E}|)=\mathcal{O}(|\mathcal{V}|\cdot d+K\cdot|\mathcal{E}|),$ (58)
dominated by feature matrices and sparse attention/distance matrices.
A.7 Stability Analysis
Theorem A.12 (Stability of DRTR).